# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v3

@@ 3 Algorithmic Approach: Subsampled Value Iteration
8:        Choose $\Delta_{i}$ uniformly at random from (${[n]\setminus i\choose k-1}$ and let $a_{i}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{i}(t),s_{\Delta_{i}}(t))]_{l}$.
@@ 4 Theoretical Guarantees and Analysis Approach
Let $\smash{\hat{Q}^{0}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta})=0}$. For $\smash{t>0}$, let $\smash{\hat{Q}_{k}^{t+1}=\hat{\mathcal{T}}_{k}\hat{Q}_{k}^{t}}$, where $\hat{\mathcal{T}}_{k}$ is defined for $k\leq n$ in Equation 11. Then, $\hat{\mathcal{T}}_{k}$ is also a $\gamma$-contraction with fixed-point $\hat{Q}_{k}^{*}$. By the law of large numbers, $\smash{\lim_{m\to\infty}\hat{\mathcal{T}}_{k,m}=\hat{\mathcal{T}}_{k}}$ and $\smash{\|\hat{Q}_{k,m}^{\mathrm{est}}-\hat{Q}_{k}^{*}\|_{\infty}\to 0}$ as $m\to\infty$. For finite $m$, $\epsilon_{k,m}\coloneq\|\hat{Q}_{k,m}^{\mathrm{est}}-\hat{Q}_{k}^{*}\|_{\infty}$ is the well-studied Bellman noise. To compare the performance between $\pi^{*}$ and ${\pi}_{k}^{\mathrm{est}}$, we define the value function of a policy $\pi$:
@@
We prove Theorem 4.4 in Appendix G, and provide a proof sketch in Appendix D. We also generalize the result to stochastic rewards in Appendix H.
To control the Bellman noise $\epsilon_{k,m}$, we show that for sufficiently many samples $m^{*}$, $\epsilon_{k,m^{*}}$ also decays on the order of $\tilde{O}(1/\sqrt{k})$ with high probability. For this, we introduce Lemma 4.3:
@@
be the number of samples in Equation 8. If the number of iterations $T$ satisfies $\smash{T\geq\frac{2}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}}$ then with probability at least $\smash{1-\frac{1}{100e^{k}}}$, the Bellman error satisfies $\smash{\epsilon_{k,m^{*}}\leq\tilde{O}(\frac{1}{\sqrt{k}})}$.
We defer the proof of Lemma 4.3 to Section G.1. To simplify notation, let ${\pi}_{k}^{\mathrm{est}}\coloneq{\pi}_{k,m^{*}}^{\mathrm{est}}$. Then, by combining Lemma 4.3 and Theorem 4.2, we obtain our main result in Theorem 4.4:
@@
Let ${\pi}^{\mathrm{est}}_{k}$ denote the learned policy from SUBSAMPLE-MFQ: Execution where the number of samples $m$ is determined in Lemma 4.3. Suppose $T\geq\frac{2}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}$. Then, $\forall s_{0}\in\mathcal{S}\coloneq\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$, with probability at least $1-1/100e^{k}$, we have  [^1] The $1/100e^{k}$ term can be replaced to any arbitrary $\delta>0$ at the cost of attaching logarithmic dependencies of $\delta$ on the $4/\sqrt{k}$ term in the error bound.,
@@
Our analysis hinges on two non-trivial technical steps. First, for an intermediary step in Theorem E.3 we establish that TV distance, rather than the stronger KL-divergence, is the correct metric to use (as the KL-divergence between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$ decays too slowly as $k\to n$, as we show in Lemma F.6). This requires exploiting a recent extension of the DKW inequality to sampling without replacement (Anand and Qu, 2024), and showing that the transition dynamics we study satisfies the data-processing inequality for TV-distance [^2] This is not true for general transition functions, since TV-distance is not an $f$-divergence.. Secondly, we adapt the celebrated performance-difference lemma (Kakade and Langford, 2002) to our multi-agent setting, which entails a principled analysis and careful probabilistic argument (to which we refer the reader to Appendix G).
@@
We also extend the formulation of SUBSAMPLE-MFQ to off-policy $Q$-learning Chen et al. (2021b); Chen and Maguluri (2022a), which replaces the generative oracle assumption with a stochastic approximation scheme that learns an approximately optimal policy using historical data. Appendix I provides theoretical guarantees with a similar decaying optimality gap as in Theorem 4.4.
@@
The asymptotic sample complexity of Algorithm 1 for learning $\hat{\pi}^{\mathrm{est}}_{k}$ for a fixed $k$ is $\min\{\tilde{O}(|\mathcal{Z}_{l}|^{k}|\mathcal{S}_{g}||\mathcal{A}_{g}|),\tilde{O}(k^{|\mathcal{Z}_{l}|}|\mathcal{Z}_{l}||\mathcal{S}_{g}||\mathcal{A}_{g}|)\}$, which is at least polynomially faster the standard $Q$-learning or mean-field value iteration as discussed in Remark 3.1. By Theorem 4.4, as $k\to n$, the optimality gap decays, revealing a fundamental trade-off in the choice of $k$: increasing $k$ improves the performance of the policy, but increases the size of the $Q$-function. We explore this trade-off further in our experiments. If we set $k=O(\log n)$, this leads to a sample complexity of $\min\{\tilde{O}(n^{\log|\mathcal{Z}_{l}|}|\mathcal{S}_{g}||\mathcal{A}_{g}|),\tilde{O}((\log n)^{|\mathcal{Z}_{l}|}|\mathcal{S}_{g}||\mathcal{A}_{g}|)\}$. This is an exponential speedup on the complexity from mean-field value iteration (from $\mathrm{poly}(n)$ to $\mathrm{poly}(\log n)$), as well as over traditional value-iteration (from $\mathrm{exp}(n)$ to $\mathrm{poly}(n)$), where the optimality gap decays to 00 with rate $O(\frac{1}{\sqrt{\log n}})$.
@@
If $k=O(\log n)$, SUBSAMPLE-MFQ handles $|\mathcal{E}|\leq O(\log n/\log\log n)$ types of local agents, since the run-time of the learning algorithm becomes $\mathrm{poly}(n)$. This surpasses the previous heterogeneity capacity from Mondal et al. (2022), which only handles constant $|\mathcal{E}|\leq O(1)$.
In the non-tabular setting with infinite state/action spaces, one could replace the $Q$-learning algorithm with any arbitrary value-based RL method that learns $\hat{Q}_{k}$ with function approximation (Sutton et al., 1999) such as deep $Q$-networks (Silver et al., 2016). Doing so raises an additional error that factors into Theorem 4.4. We formalize this below.
@@
We defer the proof of Theorem 4.10 to Appendix J.
@@ 5 Conclusion and Future Works
Our current work assumes that the global and local agents cooperate to optimize a structured reward under a specific dynamic model. While this model is more general than the federated learning setting, one direction would be to extend our algorithms and analysis to weaker network assumptions. We believe our framework, which can handle dense subgraphs, as well as expander-graph decompositions (Anand and Umans, 2023) may be amenable for this. Secondly, our current work incorporates mild heterogeneity among agents and assumes they are cooperative; thus, another avenue would be to consider settings with competitive agents or more complex agent heterogeneity. Finally, it would be exciting to generalize this work to the online no-regret setting.
@@ Acknowledgements.
We gratefully acknowledge insightful discussions with Siva Theja Maguluri, Yiheng Lin, Yi Wu, and Sarah Liaw.
@@ Appendix A Mathematical Background and Additional Remarks
  • Appendix B presents numerical simulations on the performance of SUBSAMPLE-MFQ on the Gaussian squeeze and constrained exploration tasks.
  • Appendix C presents notation and basic lemmas involving the learned $\hat{Q}_{k}$-function, as well as a stable and practical implementation of Algorithm 1.
  • Appendix D presents a proof sketch of our main result in Theorem 4.4.
  • Appendix E presents the proof of the Lipschitz continuity between $\hat{Q}_{k}$ and $Q^{*}$
  • Appendix F presents our bound on the TV-distance in Theorem D.2.
  • Appendix G proves the bound on the optimality gap between the learned policy $\tilde{\pi}_{k,m}^{\mathrm{est}}$ and the optimal policy $\pi^{*}$.
  • Appendix H presents an extension of the result to stochastic rewards.
  • Appendix I presents an extension of the result to off-policy learning that follows by using historical data.
  • Appendix J presents an extension of the result to continuous state/action spaces under a linear MDP assumption.
@@
(${[m]\choose k}$ The set of $k$-sized subsets of $\{1,\dots,m\}$;
@@ Appendix B Numerical Experiments
This section provides numerical simulations for the examples outlined in Section 2. All experiments were run on a 2-core CPU server with 12GB RAM. We chose a parameter complexity for each simulation that was sufficient to emphasize characteristics of the theory, such as the complexity improvement and the decaying optimality gap.
Figure 2: a) Reward optimality gap (log scale) with ${\pi}_{k,m}^{\mathrm{est}}$ running $300$ iterations. b) Computation time (in minutes) against sampling parameter $k$, for $k\leq n=8$, to learn policy $\hat{\pi}_{k,m}^{\mathrm{est}}$. c) Discounted cumulative rewards for $k\leq n=50$.
@@ B.1 Constrained Exploration
[…]
@@
For this task, we ran a simulation with $n=8$ agents, with $m=20$ samples in the empirically adapted Bellman operator. We provide simulation results in Figure 2a. We observe monotonic improvements in the cumulative discounted rewards as $k\to n$. Since $k=n$ recovers value-iteration and mean-field MARL algorithms, the reward at $k=n$ is the baseline we compare our algorithm to. When $k<n$, we observe that the reward accrued by SUBSAMPLE-MFQ is only marginally less than the reward gained by value-iteration.
@@ B.2 Gaussian Squeeze (GS)
For this task, we ran a small-scale simulation with $n=8$ agents, and a large-scale simulation with $n=50$ agents, and used $m=20$ samples in the empirical Bellman operator. We provide simulation results in Figure 2b and Figure 2c, where Figure 2b demonstrates the exponential improvement in computational complexity of SUBSAMPLE-MFQ, and Figure 2c demonstrates a monotonic improvement in the cumulative rewards, consistent with Theorem 4.4. Here, both metrics outperform the mean-field value iteration benchmark.

### v4

@@ 3 Algorithmic Approach: Subsampled Value Iteration
8:   Choose $\Delta_{i}$ uniformly at random from $n]\setminus i\choose k-1$} and let $a_{i}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{i}(t),s_{\Delta_{i}}(t))]_{l}$.
@@ 4 Theoretical Guarantees and Analysis Approach
Let $\smash{\hat{Q}^{0}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta})=0}$. For $\smash{t>0}$, let $\smash{\hat{Q}_{k}^{t+1}=\hat{\mathcal{T}}_{k}\hat{Q}_{k}^{t}}$, where $\hat{\mathcal{T}}_{k}$ is defined for $k\leq n$ in Equation˜13. Then, $\hat{\mathcal{T}}_{k}$ is also a $\gamma$-contraction with fixed-point $\hat{Q}_{k}^{*}$. By the law of large numbers, $\smash{\lim_{m\to\infty}\hat{\mathcal{T}}_{k,m}=\hat{\mathcal{T}}_{k}}$ and $\smash{\|\hat{Q}_{k,m}^{\mathrm{est}}-\hat{Q}_{k}^{*}\|_{\infty}\to 0}$ as $m\to\infty$. For finite $m$, $\epsilon_{k,m}\coloneq\|\hat{Q}_{k,m}^{\mathrm{est}}-\hat{Q}_{k}^{*}\|_{\infty}$ is the well-studied Bellman noise. To compare the performance between $\pi^{*}$ and ${\pi}_{k}^{\mathrm{est}}$, we define the value function of a policy $\pi$:
@@
We prove Theorem˜4.4 in Appendix˜G, and provide a proof sketch in Appendix˜D. We also generalize the result to stochastic rewards in Appendix˜H.
To control the Bellman noise $\epsilon_{k,m}$, we show that for sufficiently many samples $m^{*}$, $\epsilon_{k,m^{*}}$ also decays on the order of $\tilde{O}(1/\sqrt{k})$ with high probability. For this, we introduce Lemma˜4.3:
@@
be the number of samples in Equation˜10. If the number of iterations $T$ satisfies $\smash{T\geq\frac{2}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}}$ then with probability at least $\smash{1-\frac{1}{100e^{k}}}$, the Bellman error satisfies $\smash{\epsilon_{k,m^{*}}\leq\tilde{O}(\frac{1}{\sqrt{k}})}$.
We defer the proof of Lemma˜4.3 to Section˜G.1. To simplify notation, let ${\pi}_{k}^{\mathrm{est}}\coloneq{\pi}_{k,m^{*}}^{\mathrm{est}}$. Then, by combining Lemma˜4.3 and Theorem˜4.2, we obtain our main result in Theorem˜4.4:
@@
Let ${\pi}^{\mathrm{est}}_{k}$ denote the learned policy from SUBSAMPLE-MFQ: Execution where the number of samples $m$ is determined in Lemma˜4.3. Suppose $T\geq\frac{2}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}$. Then, $\forall s_{0}\in\mathcal{S}\coloneq\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$, with probability at least $1-1/100e^{k}$, we have  [^2] The $1/100e^{k}$ term can be replaced to any arbitrary $\delta>0$ at the cost of attaching logarithmic dependencies of $\delta$ on the $4/\sqrt{k}$ term in the error bound.,
@@
One could also derive an alternate bound that retains the $T$ factor via a $\gamma^{T}$ dependence in the final bound (which shows an exponentially decaying error with the time horizon $T$). Moreover, in Lemma˜G.10, we show that the query complexity is on the order of $O(mT|S_{g}||S_{l}|^{k}|A_{g}||A_{l}|^{k})$, and we bound $T$ and set $m=\mathrm{poly}(1/(1-\gamma))$ to attain the $1/\sqrt{k}$ rate.
@@
Additionally, our $\mathrm{poly}(1/(1-\gamma))$-dependence may be loose since we do not use more complicated variance reduction techniques as in Sidford et al. (2018a, b); Wainwright (2019); Jin et al. (2024) to optimize the number of samples $m$ which is used to bound the Bellman error $\epsilon_{k,m}$. Moreover, incorporating variance reduction would significantly complicate the algorithm and intuition.
Our analysis hinges on two non‐trivial technical steps. Firstly, for an intermediary step in Theorem˜E.3 we establish that TV distance, rather than the stronger KL-divergence, is the correct metric to use (as the KL-divergence between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$ decays too slowly as $k\to n$, as we show in Lemma˜F.6). This requires exploiting a recent extension of the DKW inequality to sampling without replacement (Anand and Qu, 2024), and showing that the transition dynamics we study saturates the data‐processing inequality for TV-distance. Secondly, we adapt the celebrated performance‐difference lemma (Kakade and Langford, 2002) to our multi-agent setting, which entails a principled analysis and careful probabilistic argument (to which we refer the reader to Appendix˜G).
@@
We also extend the formulation of SUBSAMPLE-MFQ to off-policy $Q$-learning Chen et al. (2021b); Chen and Maguluri (2022a), which replaces the generative oracle assumption with a stochastic approximation scheme that learns an approximately optimal policy using historical data. Appendix˜I provides theoretical guarantees with a similar decaying optimality gap as in Theorem˜4.4.
Remark 4.8.
The asymptotic sample complexity of Algorithm˜1 for learning $\hat{\pi}^{\mathrm{est}}_{k}$ for a fixed $k$ is $\min\{\tilde{O}(|\mathcal{Z}_{l}|^{k}|\mathcal{S}_{g}||\mathcal{A}_{g}|),\tilde{O}(k^{|\mathcal{Z}_{l}|}|\mathcal{Z}_{l}||\mathcal{S}_{g}||\mathcal{A}_{g}|)\}$, which is at least polynomially faster the standard $Q$-learning or mean-field value iteration as discussed in Remark˜3.1. By Theorem˜4.4, as $k\to n$, the optimality gap decays, revealing a fundamental trade-off in the choice of $k$: increasing $k$ improves the performance of the policy, but increases the size of the $Q$-function. We explore this trade-off further in our experiments. If we set $k=O(\log n)$, this leads to a sample complexity of $\min\{\tilde{O}(n^{\log|\mathcal{Z}_{l}|}|\mathcal{S}_{g}||\mathcal{A}_{g}|),\tilde{O}((\log n)^{|\mathcal{Z}_{l}|}|\mathcal{S}_{g}||\mathcal{A}_{g}|)\}$. This is an exponential speedup on the complexity from mean-field value iteration (from $\mathrm{poly}(n)$ to $\mathrm{poly}(\log n)$), as well as over traditional value-iteration (from $\mathrm{exp}(n)$ to $\mathrm{poly}(n)$), where the optimality gap decays to 0 with rate $O(\frac{1}{\sqrt{\log n}})$.
Remark 4.9.
If $k=O(\log n)$, SUBSAMPLE-MFQ handles $|\mathcal{E}|\leq O(\log n/\log\log n)$ types of local agents, since the run-time of the learning algorithm becomes $\mathrm{poly}(n)$. This surpasses the previous heterogeneity capacity from Mondal et al. (2022), which only handles constant $|\mathcal{E}|\leq O(1)$. Increasing the agent heterogeneity using this type formulation does make the algorithm somewhat more expensive since it factors into the query complexity via the state space of the local agents; however, since the optimality gap of the learned policy is on the order of $\tilde{O}(1/\sqrt{k})$ (modulo very small $\sqrt{\log|\mathcal{S}_{l}||\mathcal{A}_{l}|}$ factors), increasing the amount of agent heterogeneity does not degrade the quality of the learned policy as the theoretical bound does not get worse. Moreover, recent methods from the graphon mean-field MARL community might be able to enable stronger heterogeneity in the system (Cui and Koeppl, 2022; Anand and Liaw, 2025; Hu et al., 2023).
In the non-tabular setting with infinite state/action spaces, one could replace the $Q$-learning algorithm with any arbitrary value-based RL method that learns $\hat{Q}_{k}$ with function approximation (Sutton et al., 1999) such as deep $Q$-networks (Silver et al., 2016). Doing so raises an additional error that factors into Theorem˜4.4. We formalize this below.
@@
We defer the proof of Theorem˜4.12 to Appendix˜J.
@@ 5 Conclusion and Future Works
Our current work assumes that the global and local agents cooperate to optimize a structured reward under a specific dynamic model. While this model is more general than the federated learning setting, one direction would be to extend our algorithms and analysis to weaker network assumptions. We believe our framework, which can handle dense subgraphs, as well as expander-graph decompositions (Reingold, 2008; Anand and Umans, 2023; Anand, 2025) may be amenable for this. Secondly, our current work incorporates mild heterogeneity among agents and assumes they are cooperative; thus, another avenue would be to consider settings with competitive agents or more complex agent heterogeneity. Finally, it would be exciting to generalize this work to the online no-regret setting.
@@ Acknowledgements.
This work was supported by NSF Grants 2154171, CAREER Award 2339112, CMU CyLab Seed Funding, and CCF 2338816. We gratefully acknowledge insightful discussions with Siva Theja Maguluri, Yiheng Lin, Jan van den Brand, Adam Wierman, Yunbum Kook, Yi Wu, Sarah Liaw, and Rishi Veerapaneni. Finally, we thank the anonymous NeurIPS reviewers for their helpful feedback.
@@ Appendix A Mathematical Background and Additional Remarks
  • Appendix˜B presents numerical simulations on the performance of SUBSAMPLE-MFQ on the Gaussian squeeze and constrained exploration tasks.
  • Appendix˜C presents notation and basic lemmas involving the learned $\hat{Q}_{k}$-function, as well as a stable and practical implementation of Algorithm˜1.
  • Appendix˜D presents a proof sketch of our main result in Theorem˜4.4.
  • Appendix˜E presents the proof of the Lipschitz continuity between $\hat{Q}_{k}$ and $Q^{*}$
  • Appendix˜F presents our bound on the TV-distance in Theorem˜D.2.
  • Appendix˜G proves the bound on the optimality gap between the learned policy $\tilde{\pi}_{k,m}^{\mathrm{est}}$ and the optimal policy $\pi^{*}$.
  • Appendix˜H presents an extension of the result to stochastic rewards.
  • Appendix˜I presents an extension of the result to off-policy learning that follows by using historical data.
  • Appendix˜J presents an extension of the result to continuous state/action spaces under a linear MDP assumption.
  • Appendix˜K presents some practical derandomized variants with randomness-sharing.
@@
$m]\choose k$} The set of $k$-sized subsets of $\{1,\dots,m\}$;
@@ Appendix B Numerical Experiments
This section provides numerical simulations for the examples outlined in Section 2. All experiments were run on a 2-core CPU server with 12GB RAM. We chose a parameter complexity for each simulation that was sufficient to emphasize characteristics of the theory, such as the complexity improvement and the decaying optimality gap. [^3] We provide supporting code for the algorithm and experiments in https://github.com/emiletimothy/Mean-Field-Subsample-Q-Learning
Figure 2: a) Reward optimality gap (log scale) with ${\pi}_{k,m}^{\mathrm{est}}$ running 300300 iterations. b) Computation time (in minutes) against sampling parameter $k$, for $k\leq n=8$, to learn policy $\hat{\pi}_{k,m}^{\mathrm{est}}$. c) Discounted cumulative rewards for $k\leq n=50$.
@@ B.1 Constrained Exploration
Let $\mathcal{S}_{g}\!=\!\mathcal{S}_{l}\!=\![6]^{2}$ where for $s_{a}\in\{\mathcal{S}_{g},\mathcal{S}_{l}\},s_{a}^{(i)}$ denotes the projection of $s_{a}$ onto its $i$’th coordinate. Further, let $\mathcal{A}_{l}=\mathcal{A}_{g}=\{\text{``up", ``down", ``right", ``left"}\}$ given formally by $\{\binom{1}{0},\binom{-1}{0},\binom{0}{1},\binom{0}{-1}\}$. Let $\Pi_{D}(x)$ denote the $\ell_{1}$-projection of $x$ onto set $D$. Then, let $s_{g}^{t+1}(s_{g}^{t},a_{g}^{t})=\Pi_{S_{g}}(s_{g}^{t}+a_{g}^{t}),s_{i}^{t+1}(s_{i}^{t},s_{g}^{t},a_{i}^{t})=\Pi_{S_{l}}(s_{i}^{t}+|s_{i}^{t}-s_{g}^{t}|+a_{i}^{t})$. We let the global agent’s reward be $r_{g}(s_{g},a_{g})=2\sum_{i=1}^{3}\mathbbm{1}\{a_{g}={1\choose 0},s_{g}^{(1)}\leq i\}+2\sum_{i=3}^{6}\mathbbm{1}\{a_{g}={-1\choose 0},s_{g}^{(1)}>i\}$, and the local agent’s rewards be $r_{l}(s_{i},s_{g},a_{i})=\mathbbm{1}\{a_{i}^{t}\neq 0\}+6\cdot\mathbbm{1}\{s_{i}^{(1)}=s_{g}^{(1)}\}+2\cdot\|s_{i}^{t}-s_{g}^{t}\|_{1}$.
@@
For this task, we ran a simulation with $n=8$ agents, with $m=20$ samples in the empirically adapted Bellman operator. We provide simulation results in Figure˜2a. We observe monotonic improvements in the cumulative discounted rewards as $k\to n$. Since $k=n$ recovers value-iteration and mean-field MARL algorithms, the reward at $k=n$ is the baseline we compare our algorithm to. When $k<n$, we observe that the reward accrued by SUBSAMPLE-MFQ is only marginally less than the reward gained by value-iteration.
@@ B.2 Gaussian Squeeze (GS)
For this task, we ran a small-scale simulation with $n=8$ agents, and a large-scale simulation with $n=50$ agents, and used $m=20$ samples in the empirical Bellman operator. We provide simulation results in Figure˜2b and Figure˜2c, where Figure˜2b demonstrates the exponential improvement in computational complexity of SUBSAMPLE-MFQ, and Figure˜2c demonstrates a monotonic improvement in the cumulative rewards, consistent with Theorem˜4.4. Here, both metrics outperform the mean-field value iteration benchmark.